OpenStudy (anonymous):

I need help with this question: An airliner flies 2000 km at a speed of 1600 km/h and then returns due to bad weather at a speed of 1000 km/h. Find the average speed for the whole trip.

OpenStudy (gorv):

average speed =total distance/total time

OpenStudy (gorv):

here total distance equal = total distance of return trip=2000+2000 time =distance / speed=2000/1600 + 2000/1000

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

Can you please explain again?

OpenStudy (anonymous):

Somebody? Anybody?

OpenStudy (phi):

First, speed* time = distance if we divide both sides by "speed" we get time = distance/speed Using 2000 km at a speed of 1600 km/h how long did it take the plane to travel 2000 km ? can you do that ?

OpenStudy (anonymous):

It's okay, I already solved it

OpenStudy (anonymous):

I'll show you what I did

OpenStudy (anonymous):

First, we must get the time for the first part of the journey. That would be: 2000/1600= 1.25 hours (multiplied by 60= 75 mins, so 1 hour 15 mins) Then, 2000/1000= 2 hours 2+1.25= 3.25 hours ( 3hours and 15 mins) So, to get the avg. speed: Total distance, which is, 4000 divided by 3.25 hours 4000/3.25=1230 km/h (about) Thanks for helping anyway : )

OpenStudy (phi):

yes

OpenStudy (phi):

though you rounded to the nearest tenth. more accurately 1230.77 km/h

OpenStudy (anonymous):

Yes, that's what I did

OpenStudy (anonymous):

I rounded it to the nearest tenth

OpenStudy (anonymous):

But it doesn't really matter, right?